OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) + 6

OpenStudy (anonymous):

3x+3 = -2x +2+6 3x +3 = -2x +8 5x= 5 x=1

OpenStudy (anonymous):

x=1